# A Generic Machine Learning Framework for Radio Frequency Fingerprinting

### v1

@@ 3 Problem Formulation
In this section, we formally introduce the framework for learning general-purpose RF fingerprint representations using ML. Let us assume that we have a dataset $\mathcal{D}=({\bf{X}},{\bf{Y}})\in\mathcal{X}\times\mathcal{Y}$ where $\bf{X}$} and $\bf{Y}$} are sampled from the unknown joint probability distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$, where $\mathcal{X}$ is the RF feature space and $\mathcal{Y}$ is the label space. We define a $\it{task}$} $\mathcal{T}$ to be composed of the dataset $\mathcal{D}$, a loss function $\mathcal{L}$ and model parameterizations ${\bf{g}}_{\phi}$ and ${\bf{F}}_{\theta}$, where ${\bf{g}}_{\phi}$ is the task head, and ${\bf{F}}_{\theta}$ is a vector-valued RFF function that contains, or is equivalent to, the RF fingerprint head ${\bf{f}}_{\theta}:\mathcal{X}\mapsto\mathcal{Z}$, where $\mathcal{Z}$ is the RFF representation space. In general, ${\bf{F}}_{\theta}:\mathcal{X}\times\cdots\times\mathcal{X}\mapsto[\mathcal{Z}]^{l}$, where $l$ denotes the number of RF fingerprint head projections that have been made and $\mathcal{Z}]^{l}\triangleq\mathbb{R}^{d\times l}$. For the majority of tasks, such as SEI, $l=1$ and the vector-valued function ${\bf{F}}_{\theta}\equiv{\bf{f}}_{\theta}$ whereas for other tasks, such as EDA, $l>1$.
@@
The composite function ${\bf{h}}_{\xi}({\bf{x}})\triangleq{\bf{g}}_{\phi}({\bf{F}}_{\theta}({\bf{x}}))$ projects the input data ${\bf{x}}\in{\bf{X}}$ to a prediction $\hat{{\bf{y}}}$, which is computed for every ${\bf{x}}\in{\bf{X}}$, forming a predictive labels set $\hat{{\bf{Y}}}=\{\hat{{\bf{y}}}|\hat{{\bf{y}}}={\bf{h}}_{\xi}({\bf{x}}),{\bf{x}}\in{\bf{X}}\}$. The loss function $\mathcal{L}$ measures the distance between any pairs of $({\bf{y}},\hat{{\bf{y}}})\in{\bf{Y}}\times\hat{{\bf{Y}}}$ and is used to supervise model training, and ${\bf{g}}_{\phi}:[\mathcal{Z}]^{l}\mapsto\mathcal{Y}$ is a mapping from $\mathcal{Z}]^{l}$ to the label space $\mathcal{Y}$.
@@
Notationally, we represent the $k$th $\it{task}$} with superscripts (e.g. $t$th task dataset $\mathcal{D}^{(t)}$), where each task is assumed to be sampled from the probability distribution of tasks $\mathcal{P}_{\mathcal{T}}$ and $\mathcal{D}^{(t)}=({\bf{X}}^{(t)},{\bf{Y}}^{(t)})=\{({\bf{x}}_{i}^{(t)},y_{i}^{(t)})\}_{i=1}^{n_{t}}\sim\mathcal{P}_{\mathcal{X}^{(t)}{\mathcal{Y}^{(t)}}}$. Therefore, we also have the $t$th task predictive labels set $\hat{{\bf{Y}}}^{(t)}$. Note that for some tasks, the parameter set $\phi$ is fixed since the function $\bf{g}$} is a deterministic function (e.g. distance measure) but for others is learnable. For example, ${\bf{g}}_{\phi}$ maps the lower-dimensional representation to the classification label space for a SEI task, whereas for an EDA task, ${\bf{g}}_{\phi}$ is a distance measure (e.g. cosine distance ) between two RF input embeddings. Hence, $\phi$ is not strictly learnable whereas $\theta$ is.
3.1 Multi Task RF Fingerprint Learning
@@
We first introduce the following optimization problem:
@@
where ${\bf{h}}_{\xi_{t}}^{J,(t)}({\bf{X}}^{(t)})={\bf{g}}_{\phi_{t}}^{(t)}\left({\bf{F}}_{\theta}^{(t)}\left({\bf{X}}^{(t)}\right)\right)$, with superscript $J$ representing that $\theta$ is jointly learnt across tasks, $\Xi=\Theta\times{\bf{\Phi}}$, $\Theta\equiv\mathbb{R}^{d},d\in\mathbb{Z}^{+}$, ${\bf{\Phi}}=\Phi_{1}\times\cdots\times\Phi_{|\mathcal{P}_{\mathcal{T}}|}$, $\Phi_{j}\equiv\mathbb{R}^{c_{j}},c_{j}\in\mathbb{Z}^{+}$, with $\theta^{*}\in\Theta$, ${\boldsymbol{\phi}}^{*}=(\phi^{*}_{1},\cdots,\phi^{*}_{|\mathcal{P}_{\mathcal{T}}|})$ with $\phi^{*}_{j}\in\Phi_{j}\;\forall j\in\{1,\cdots,|\mathcal{P}_{\mathcal{T}}|\}$. Upon solving the above optimization task, we jointly learn the RF fingerprint parameters $\theta$ and the task-specific parameter sets $\phi_{t}$, for all $t$.
We define the parameters $\xi^{*}$ as an argument minimizer of the optimization problem above. Intuitively speaking, it is expected that the resulting RF fingerprint head ${\bf{f}}_{\theta^{*}}$ has an increasingly richer RFF representation when jointly learning $\theta$ across an increasing amount of tasks that require RF fingerprint knowledge. It is assumed that the functional form of the RF fingerprint head is chosen to be the same across all tasks. In other words, ${\bf{F}}_{\theta}^{(t)}={\bf{F}}_{\theta}$ where ${\bf{F}}_{\theta}$ is chosen appropriately.
Alternatively, we can learn across all tasks but have independent parameter sets $\theta_{t}$ for each task $t$ and aggregate them post optimization. Mathematically speaking,
@@
where ${\bf{h}}_{\xi_{t}}^{I,(t)}({\bf{X}}^{(t)})={\bf{g}}_{\phi_{t}}^{(t)}\left({\bf{F}}_{\theta_{t}}^{(t)}\left({\bf{X}}^{(t)}\right)\right)$, with $I$ denoting that an independent parameter set $\theta_{t}$ is being learnt for each task. An aggregation function $a:\Theta_{1}\times\cdots\times\Theta_{\lvert\mathcal{P}_{\mathcal{T}}\rvert}\mapsto\mathbb{R}^{d}$ can then be used to combine the parameters into one parameter set $\theta^{*}$. In other words, $\theta^{*}=a({\boldsymbol{\theta}}^{*}).$ One possible composition of the aggregation function $a$ could be $a({\bf{x}})=\frac{1}{n}\sum_{i=1}^{n}x_{i}$, for example.
@@
In the discrete setting, we assume that $m$ tasks exists, where the $t$th task has a dataset $\mathcal{D}^{(t)}$, with each data point ${\bf{z}}\in\mathcal{D}^{(t)}$ being Independent Identically Distributed (IID) and assumed to be drawn from an unknown joint probability distribution $\mathcal{P}_{\mathcal{X}^{(t)}\mathcal{Y}^{(t)}}$. This results in the following optimization problems:
$\boldsymbol{\xi}_{S,J}^{*}=\mbox{arg}\min_{\xi\in\Xi}\sum_{t=1}^{m}\frac{\alpha_{t}}{n_{t}}\sum_{j=1}^{n_{t}}\mathcal{L}^{(t)}\left({\bf{h}}^{J,(t)}_{\xi_{t}}\left({\bf{x}}^{(t)}_{j}\right),{\bf{y}}^{(t)}_{j}\right)$,
$\boldsymbol{\xi}_{S,I}^{*}=\mbox{arg}\min_{\xi\in\Xi}\sum_{t=1}^{m}\frac{\alpha_{t}}{n_{t}}\sum_{j=1}^{n_{t}}\mathcal{L}^{(t)}\left({\bf{h}}^{I,(t)}_{\xi_{t}}\left({\bf{x}}^{(t)}_{j}\right),{\bf{y}}^{(t)}_{j}\right)$,
@@
where the argument minimizers ${\boldsymbol{\xi}}_{S,J}^{*}=(\boldsymbol{\phi}_{S,J}^{*},{\theta}_{S,J}^{*})=(\phi_{1}^{*},\cdots,\phi^{*}_{n},\theta^{*})$ and $\boldsymbol{\xi}_{S,I}^{*}=({\boldsymbol{\phi}}_{S,I}^{*},\boldsymbol{\theta}_{S,I}^{*})=(\phi_{1}^{*},\cdots,\phi^{*}_{m},\theta^{*}_{1},\cdots,\theta^{*}_{m})$ for the joint and individual (parameter set for each task) optimization tasks respectively, and $\alpha_{t}$ denotes the weighting of the $t$th task with $\sum_{t=1}^{n}\alpha_{t}=1$. A standard choice for this is equal weighting across tasks (e.g. $\alpha_{t}=1/n$ for all $t$), though this formulation allows for this to be selected according to preference. In the same manner as previously discussed, we can aggregate to find $\theta^{*}=a({\boldsymbol{\theta}}^{*})=a(\theta_{1}^{*},\cdots,\theta_{m}^{*})$.
@@
The resulting parameters found from solving either of these optimization problems are a function of the RFF-based tasks, data quality, loss function and model architecture. In practice, whilst $\mathcal{P}_{\mathcal{T}}$ is unknown, it is assumed that common tasks such as SEI, EDA, RF Emitter Clustering (RFEC), amongst others are the modes of this distribution and approximate the expectation appropriately. Therefore, it is likely that approximate solutions to this optimization problem can be found by jointly considering many popular RF fingerprint-based tasks.
@@
The joint optimization problems introduced here are analogous to multi-modal joint embedding learning techniques [57, 43, 2], though here the intuition is on solving different tasks in the multi-task learning (MTL) sense [49, 69, 5], rather than enriching the RF representation via multi-modal data. Both formulations are multi-objective optimization problems that serve multi-task learning [49] and are a generic approach towards learning RF fingerprint representation spaces via data-driven methods.
4 Single Task RF Fingerprint Learning
In this section, we provide descriptions of popular RF fingerprint-based tasks using the framework introduced in Section 3. It is noted that only one RFF-based task is considered at a time.
4.1 Closed Set Specific Emitter Identification
@@
The first task we define is closed set SEI (CS-SEI). In this task, the objective is to learn an accurate mapping from emitter observations (typically I/Q samples) to the correct emitter label. Assuming that $\mathcal{P}_{{\mathcal{X}}^{(i)}{\mathcal{Y}}^{(i)}}$ is the joint probability distribution of the SEI feature space $\mathcal{X}^{(i)}$ and the emitter label space $\mathcal{Y}^{(i)}$, where the superscript $i$ denotes identification task, we pose the following optimization problem to learn a mapping ${\bf{h}}_{\xi}$ from ${\mathcal{X}}^{(i)}$ to ${\mathcal{Y}}^{(i)}$:
@@
This translates as finding a suitable parameter set pair $\xi^{*}=(\phi^{*},\theta^{*})$ such that ${\bf{h}}_{\xi^{*}}({\bf{x}})={\bf{g}}_{\phi^{*}}({\bf{F}}_{\theta^{*}}({\bf{x}}))\approx y\;\forall({\bf{x}},y)\in\mathcal{X}^{(i)}\times\mathcal{Y}^{(i)}$, noting here that for CS-SEI ${\bf{F}}_{\theta^{*}}\equiv{\bf{f}}_{\theta^{*}}$.
Since the joint probability distribution is unknown, we cannot solve this optimization problem in its exact form. Instead, we minimize the empirical error, replacing the joint distribution with the sampled empirical data. We assume a dataset $\mathcal{D}^{(i)}=({\bf{X}}^{(i)},{\bf{Y}}^{(i)})$ is sampled from $\mathcal{P}_{{\mathcal{X}}^{(i)}{\mathcal{Y}}^{(i)}}$, which we expand as $\mathcal{D}^{(i)}=\{({\bf{x}}^{(i)}_{k},y^{(i)}_{k})\}_{k=1}^{n_{i}}$ where ${\bf{x}}^{(i)}_{k}\in\mathbb{R}^{d^{(i)}}$, $y_{k}\in\{1,2,\cdots,\mathcal{V}\}$, $n_{i}$ is the number of samples for the identification task and $\mathcal{V}$ is the number of unique emitters. We assume that each data point ${\bf{z}}=({\bf{x}},y)\in\mathcal{D}^{(i)}$ is sampled IID from $\mathcal{P}_{{\mathcal{X}}^{(i)}{\mathcal{Y}}^{(i)}}$. Then, the CS-SEI optimization problem is defined as:
@@
4.2 RFF-based Emitter Data Association
The objective in this task is to learn a model that can determine whether a pair of emitter observations originate from the same emitter or not. In other words, we are interested in finding a function ${\bf{h}}_{\xi}:{\mathcal{X}}^{(a)}\mapsto\mathcal{Y}^{(a)}=\{0,1\}$, where ${\mathcal{X}}^{(a)}$ is the EDA feature space, and labels 0 and 1 denotes no match and match respectively. In this case, the model prediction $\hat{y}={\bf{h}}^{(a)}_{\xi}({\bf{x}})$ is parameterized as follows:

### v2

@@ III Problem Statement and RFF Framework
We formally introduce here the framework for learning general-purpose RF fingerprint representations using ML. Let us assume that we have a dataset $\mathcal{D}=({\bf{X}},{\bf{Y}})\in\mathcal{X}\times\mathcal{Y}$ where $\bf{X}$} and $\bf{Y}$} are sampled from the unknown joint probability distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$, where $\mathcal{X}$ is the RF feature space and $\mathcal{Y}$ is the label space. We define a $\it{task}$} $\mathcal{T}$ to be composed of the dataset $\mathcal{D}$, a loss function $\mathcal{L}$ and model parameterisations of ${\bf{g}}_{\phi}$ and ${\bf{F}}_{\theta}$, where ${\bf{g}}_{\phi}$ is the task head, and ${\bf{F}}_{\theta}$ is a vector-valued RFF function that contains, or is equivalent to, the RF fingerprint head ${\bf{f}}_{\theta}:\mathcal{X}\mapsto\mathcal{Z}$, where $\mathcal{Z}$ is the RFF representation (i.e. embeddings) space. The latter function performs the representation learning as depicted in Figure 1 (i.e. produces the sought RF fingerprint from a given input RF data). In general, ${\bf{F}}_{\theta}:\mathcal{X}\times\cdots\times\mathcal{X}\mapsto[\mathcal{Z}]^{l}$, where $l$ denotes the number of RF fingerprint head projections that have been made and $\mathcal{Z}]^{l}\triangleq\mathbb{R}^{d\times l}$. For the majority of tasks, such as SEI, $l=1$ and the vector-valued function ${\bf{F}}_{\theta}\equiv{\bf{f}}_{\theta}$ whereas for other tasks, e.g. EDA, $l>1$. Without loss of generality, ${\bf{g}}_{\phi}$ and ${\bf{F}}_{\theta}$, jointly or individually, can be ML (DNN) models with the associated (hyper)parameters (i.e. $\phi$ and $\theta$) depending on the pursued downstream task(s).
@@
The composite function ${\bf{h}}_{\xi}({\bf{x}})\triangleq{\bf{g}}_{\phi}({\bf{F}}_{\theta}({\bf{x}}))$ projects the input data ${\bf{x}}\in{\bf{X}}$ to a prediction $\hat{{\bf{y}}}$, which is computed for every ${\bf{x}}\in{\bf{X}}$, forming a predictive labels set $\hat{{\bf{Y}}}=\{\hat{{\bf{y}}}|\hat{{\bf{y}}}={\bf{h}}_{\xi}({\bf{x}}),{\bf{x}}\in{\bf{X}}\}$. The loss function $\mathcal{L}$ measures the distance between any pairs of $({\bf{y}},\hat{{\bf{y}}})\in{\bf{Y}}\times\hat{{\bf{Y}}}$ and is used to supervise the model training, and ${\bf{g}}_{\phi}:[\mathcal{Z}]^{l}\mapsto\mathcal{Y}$ is a mapping from $\mathcal{Z}]^{l}$ to the label space $\mathcal{Y}$, where relevant.
@@
Notationally, we represent the $k$th $\it{task}$} with superscripts (e.g. $t$th task dataset $\mathcal{D}^{(t)}$), where each task is assumed to be sampled from the probability distribution of tasks $\mathcal{P}_{\mathcal{T}}$ and $\mathcal{D}^{(t)}=({\bf{X}}^{(t)},{\bf{Y}}^{(t)})=\{({\bf{x}}_{i}^{(t)},y_{i}^{(t)})\}_{i=1}^{n_{t}}\sim\mathcal{P}_{\mathcal{X}^{(t)}{\mathcal{Y}^{(t)}}}$. Therefore, we also have the $t$th task predictive labels set $\hat{{\bf{Y}}}^{(t)}$. Note that for some tasks, the parameter set $\phi$ is fixed since the function $\bf{g}$} is a deterministic function (e.g. distance measure) but for others is learnable. For example, ${\bf{g}}_{\phi}$ maps the inferred lower-dimensional characteristic representation (i.e. RF fingerprint) to the classification label space for a SEI task, whereas for an EDA task, ${\bf{g}}_{\phi}$ is a distance measure (e.g. cosine distance) between two RF input embeddings. Hence, $\phi$ is not strictly learnable whereas $\theta$ (i.e. the representation learning model parameters such TAas the weights in a DNN) is.
@@
Within a multi-task RFF context, we first introduce the following optimisation problem:
@@
where ${\bf{h}}_{\xi_{t}}^{J,(t)}({\bf{X}}^{(t)})={\bf{g}}_{\phi_{t}}^{(t)}\left({\bf{F}}_{\theta}^{(t)}\left({\bf{X}}^{(t)}\right)\right)$, with superscript $J$ for $\theta$ that is jointly learnt across tasks, $\Xi=\Theta\times{\bf{\Phi}}$, $\Theta\equiv\mathbb{R}^{d},d\in\mathbb{Z}^{+}$, ${\bf{\Phi}}=\Phi_{1}\times\cdots\times\Phi_{|\mathcal{P}_{\mathcal{T}}|}$, $\Phi_{j}\equiv\mathbb{R}^{c_{j}},c_{j}\in\mathbb{Z}^{+}$, with $\theta^{*}\in\Theta$, ${\bm{\phi}}^{*}=(\phi^{*}_{1},\cdots,\phi^{*}_{|\mathcal{P}_{\mathcal{T}}|})$ with $\phi^{*}_{j}\in\Phi_{j}\;\forall j\in\{1,\cdots,|\mathcal{P}_{\mathcal{T}}|\}$. Upon solving the above optimisation task, we jointly learn the RF fingerprinting model parameters $\theta$ and the task-specific parameters $\phi_{t}$, for all $t$.
We define the parameters $\xi^{*}$ as an argument minimiser of the optimisation problem above. Intuitively speaking, it is expected that the resulting RFF head ${\bf{f}}_{\theta^{*}}$ has an increasingly richer fingerprint representation when jointly learning $\theta$ across an increasing amount of tasks that require RF fingerprint knowledge. It is assumed that the functional form of the RF fingerprint head is chosen to be the same across all tasks. In other words, ${\bf{F}}_{\theta}^{(t)}={\bf{F}}_{\theta}$ where ${\bf{F}}_{\theta}$ is chosen appropriately.
Alternatively, we can learn across all tasks but have independent parameter sets $\theta_{t}$ for each task $t$ and aggregate them post optimisation. Mathematically speaking,
@@
where ${\bf{h}}_{\xi_{t}}^{I,(t)}({\bf{X}}^{(t)})={\bf{g}}_{\phi_{t}}^{(t)}\left({\bf{F}}_{\theta_{t}}^{(t)}\left({\bf{X}}^{(t)}\right)\right)$, with $I$ denoting that an independent parameter set $\theta_{t}$ is being learnt for each task. An aggregation function $a:\Theta_{1}\times\cdots\times\Theta_{\lvert\mathcal{P}_{\mathcal{T}}\rvert}\mapsto\mathbb{R}^{d}$ can then be used to combine the parameters into one parameter set $\theta^{*}$. In other words, $\theta^{*}=a({\bm{\theta}}^{*}).$ One possible composition of the aggregation function $a$ could be $a({\bf{x}})=\frac{1}{n}\sum_{i=1}^{n}x_{i}$, for example.
@@
In the discrete setting, we assume that $m$ tasks exists, where the $t$th task has a dataset $\mathcal{D}^{(t)}$, with each data point ${\bf{z}}\in\mathcal{D}^{(t)}$ being independent identically distributed (IID) and assumed to be drawn from an unknown joint probability distribution $\mathcal{P}_{\mathcal{X}^{(t)}\mathcal{Y}^{(t)}}$. This results in the following optimisation problems:
$\bm{\xi}_{S,J}^{*}=\mbox{arg}\min_{\xi\in\Xi}\sum_{t=1}^{m}\frac{\alpha_{t}}{n_{t}}\sum_{j=1}^{n_{t}}\mathcal{L}^{(t)}\left({\bf{h}}^{J,(t)}_{\xi_{t}}\left({\bf{x}}^{(t)}_{j}\right),{\bf{y}}^{(t)}_{j}\right)$,
$\bm{\xi}_{S,I}^{*}=\mbox{arg}\min_{\xi\in\Xi}\sum_{t=1}^{m}\frac{\alpha_{t}}{n_{t}}\sum_{j=1}^{n_{t}}\mathcal{L}^{(t)}\left({\bf{h}}^{I,(t)}_{\xi_{t}}\left({\bf{x}}^{(t)}_{j}\right),{\bf{y}}^{(t)}_{j}\right)$,
@@
where the argument minimisers ${\bm{\xi}}_{S,J}^{*}=(\bm{\phi}_{S,J}^{*},{\theta}_{S,J}^{*})=(\phi_{1}^{*},\cdots,\phi^{*}_{n},\theta^{*})$ and $\bm{\xi}_{S,I}^{*}=({\bm{\phi}}_{S,I}^{*},\bm{\theta}_{S,I}^{*})=(\phi_{1}^{*},\cdots,\phi^{*}_{m},\theta^{*}_{1},\cdots,\theta^{*}_{m})$ for the joint and individual (parameter set for each task) optimisation tasks respectively, and $\alpha_{t}$ denotes the weighting of the $t$th task with $\sum_{t=1}^{n}\alpha_{t}=1$. A standard choice for this is equal weighting across tasks (e.g. $\alpha_{t}=1/n$ for all $t$), though this formulation allows for this to be selected according to preference. In the same manner as previously discussed, we can perform aggregation to find $\theta^{*}=a({\bm{\theta}}^{*})=a(\theta_{1}^{*},\cdots,\theta_{m}^{*})$.
@@
The resulting parameters found from solving either of these optimisation problems are a function of the RFF-based tasks, data quality, loss function and model architecture. In practice, whilst $\mathcal{P}_{\mathcal{T}}$ is unknown, it is assumed that common tasks such as SEI, EDA and RFEC, amongst others, are the modes of this distribution and approximate the expectation appropriately. Therefore, it is likely that approximate solutions to this optimisation problem can be found by jointly considering many popular RF fingerprint-based tasks.
@@
The joint optimisation problems introduced here are analogous to multi-modal joint embedding learning techniques [56, 57, 58], though here the intuition is on solving different tasks in the multi-task learning (MTL) sense [59, 60, 61], rather than enriching the RF representation via multi-modal data. Both formulations are multi-objective optimisation problems that serve multi-task learning [59] and are a generic approach towards learning RF fingerprint representation spaces via data-driven methods.
IV Single Task RF Fingerprint Learning
We now present a few popular RF fingerprint-based tasks, each considered individually.
IV-A Closed-Set Specific Emitter Identification
@@
The first task we define is closed-set SEI (CS-SEI). The objective is to learn an accurate mapping from emitter observations (e.g. I/Q samples) to the correct emitter label. Assuming that $\mathcal{P}_{{\mathcal{X}}^{(i)}{\mathcal{Y}}^{(i)}}$ is the joint probability distribution of the SEI feature space $\mathcal{X}^{(i)}$ and the emitter label space $\mathcal{Y}^{(i)}$, where the superscript $i$ denotes identification task. We pose the following optimisation problem to learn a mapping ${\bf{h}}_{\xi}$ from ${\mathcal{X}}^{(i)}$ to ${\mathcal{Y}}^{(i)}$:
@@
This translates to finding a suitable parameter set pair $\xi^{*}=(\phi^{*},\theta^{*})$ such that ${\bf{h}}_{\xi^{*}}({\bf{x}})={\bf{g}}_{\phi^{*}}({\bf{F}}_{\theta^{*}}({\bf{x}}))\approx y\;\forall({\bf{x}},y)\in\mathcal{X}^{(i)}\times\mathcal{Y}^{(i)}$, noting here that for CS-SEI ${\bf{F}}_{\theta^{*}}\equiv{\bf{f}}_{\theta^{*}}$.
Since the joint probability distribution is unknown in practice, we cannot directly solve this optimisation problem in its exact form. Instead, we minimise the empirical error, replacing the joint distribution with the sampled empirical data. We assume a dataset $\mathcal{D}^{(i)}=({\bf{X}}^{(i)},{\bf{Y}}^{(i)})$ is sampled from $\mathcal{P}_{{\mathcal{X}}^{(i)}{\mathcal{Y}}^{(i)}}$, which we expand as $\mathcal{D}^{(i)}=\{({\bf{x}}^{(i)}_{k},y^{(i)}_{k})\}_{k=1}^{n_{i}}$ where ${\bf{x}}^{(i)}_{k}\in\mathbb{R}^{d^{(i)}}$, $y_{k}\in\{1,2,\cdots,\mathcal{V}\}$, $n_{i}$ is the number of samples for the identification task and $\mathcal{V}$ is the number of unique emitters in the considered closed-set. We assume that each data point ${\bf{z}}=({\bf{x}},y)\in\mathcal{D}^{(i)}$ is IID sampled from $\mathcal{P}_{{\mathcal{X}}^{(i)}{\mathcal{Y}}^{(i)}}$. The CS-SEI optimisation problem is defined as:
@@
IV-B RFF-based Emitter Data Association
The objective in this task is to learn a model that can determine whether a pair of emitter observations originate from the same emitter or not. In other words, we are interested in finding a function ${\bf{h}}_{\xi}:{\mathcal{X}}^{(a)}\mapsto\mathcal{Y}^{(a)}=\{0,1\}$, where ${\mathcal{X}}^{(a)}$ is the EDA feature space, and labels 0 and 1 denotes no match and match, respectively. In this case, the model prediction $\hat{y}={\bf{h}}^{(a)}_{\xi}({\bf{x}})$ is parameterised as follows: